OpenStudy (anonymous):

Let y = 4*sqrtx Estimate the difference delta y when x = 5 using a linear approximation , given that delta x = 0.3

OpenStudy (bahrom7893):

@TuringTest I forgot my linear approximations.

OpenStudy (turingtest):

This is phrased strangely for me, but I assume they want you to do is say for \(x\approx x_0=5\) we have that\[y(x)\approx y(x_0)+y'(x_0)(x-x_0)\implies y(x)-y(x_0)\approx y'(x_0)(x-x_0)\]\[\Delta y=y(x)-y(x_0),~~~~\Delta x=x-x_0=0.3\]therefor\[\Delta y\approx y'(5)\Delta x\]

OpenStudy (turingtest):

@sam0102493 let me know if that makes sense to you